Find the remaining trigonometric ratios of θ based on the given information.
sec θ = a and θ terminates in QI

Answer by lwsshak3(11628)

Find the remaining trigonometric ratios of θ based on the given information.
sec θ = a and θ terminates in QI
***
secx=a
cosx=1/a
sinx=√1-cos^2(x)=√(1-1/a^2)=√((a^2-1)/a^2)=√(a^2-1)/a
cscx=a/√(a^2-1)
tanx=√(a^2-1)
cotx=1/√(a^2-1)